OpenStudy (anonymous):

the sum of two numbers is 70. one number is 8 more than the other.whats the smaller number?

OpenStudy (anonymous):

hi drizzy

OpenStudy (anonymous):

this is a question with two unknowns , and so you need two equations to solve it

OpenStudy (anonymous):

that's the tricky part in general - figuring out the two equations from a word problem

OpenStudy (anonymous):

so here are the equations I get: x+y = 70 x = 7y

OpenStudy (anonymous):

sorry x=8y for the second

OpenStudy (anonymous):

the next step is to substitute the value for a variable in one equation into the other

OpenStudy (anonymous):

this is what I get, Isnt is x for first number, x+8 for second number, then x + (x+8) = 70, 2x +8 = 70 2x=62, x= 31

OpenStudy (anonymous):

31 + (31+8) = 31 + 39 = 70